Pelamed Pharmaceuticals has EBIT of $418 million in 2012. In​ addition, Pelamed has interest expenses of $187 million and a corporate tax rate of 30%.

a. What is​ Pelamed's 2012 net​ income?

b. What is the total of​ Pelamed's 2012 net income plus interest​ payments?

c. If Pelamed had no interest​ expenses, what would its 2012 net income​ be? How does it compare to your answer in part ​(b​)?

d. What is the amount of​ Pelamed's interest tax shield in​ 2012?
Answer #1

1.
Net Income=(EBIT-Interest)*(1-tax rate)
=(418-187)*(1-30%)
=161.70

2.
=Net Income+Interest
=(418-187)*(1-30%)+187
=348.70

3.
=(EBIT-0)*(1-tax rate)
=(418)*(1-30%)
=292.60

Lower by interest*tax rate

4.
=Interest*tax rate
=187*30%
=56.10